OpenStudy (agentjamesbond007):

if k>0, lim as x approaches k (√x-√k)/(k-x), what is the limit?

OpenStudy (anonymous):

x is the variable , k is a constant. So if we were to sub x=k into the expression we would get 0/0 (indeterminate form). So use L Hopital's rule. Differentiate top and bottom with respect to the VARIABLE (i.e. k is treated as a constant). = l\[\lim_{x \rightarrow k} \frac{ \frac{1}{2 \sqrt{x}} - 0 }{0-1} = -\frac{1}{2\sqrt{k}}\]